# Active Learning with Variational Quantum Circuits for Quantum Process Tomography

### v1

@@ Abstract
Quantum process tomography (QPT), used for reconstruction of an unknown quantum process from measurement data, is a fundamental tool for the diagnostic and full characterization of quantum systems. It relies on querying a set of quantum states as input to the quantum process. Previous works commonly use a straightforward strategy to select a set of quantum states randomly, overlooking differences in informativeness among quantum states. Since querying the quantum system requires multiple experiments that can be prohibitively costly, it is always the case that there are not enough quantum states for high-quality reconstruction. In this paper, we propose a general framework for active learning (AL) to adaptively select a set of informative quantum states that improves the reconstruction most efficiently. In particular, we introduce a learning framework that leverages the widely-used variational quantum circuits (VQCs) to perform the QPT task and integrate our AL algorithms into the query step. We design and evaluate three various types of AL algorithms: committee-based, uncertainty-based, and diversity-based, each exhibiting distinct advantages in terms of performance and computational cost. Additionally, we provide a guideline for selecting algorithms suitable for different scenarios. Numerical results demonstrate that our algorithms achieve significantly improved reconstruction compared to the baseline method that selects a set of quantum states randomly. Moreover, these results suggest that active learning based approaches are applicable to other complicated learning tasks in large-scale quantum information processing.
Corresponding author.
Introduction
With recent advances in quantum computing and machine learning, their intersection has emerged as an increasingly attractive area of research (Biamonte et al. 2017; Schuld, Sinayskiy, and Petruccione 2015). On one hand, quantum computing enhances the efficiency of machine learning algorithms by exploiting quantum advantage (Dunjko, Taylor, and Briegel 2016; Wiebe, Kapoor, and Svore 2014). On the other hand, machine learning methodologies and techniques are employed to help solving problems in quantum computation and quantum information (Carleo and Troyer 2017; Mitarai et al. 2018). In this paper, we apply active learning (AL) to solve a challenging problem of quantum process tomography (QPT).
@@ Introduction
QPT, used for reconstruction of an unknown quantum process from measurement data, is a fundamental tool for the diagnostic and full characterization of quantum systems (Chuang and Nielsen 1997). Standard quantum process tomography (SQPT) relies on an informationally complete set of quantum states as input, which inevitably leads to an algorithmic complexity in terms of the number of measurements and classical post-processing that scales exponentially with the system size (D’Ariano and Presti 2001). Due to these limitations, SQPT has only been experimentally implemented up to three qubits (Krinner et al. 2020; Muñoz-Gil et al. 2020). However, with the rapid advancement of quantum computing hardware, there is a growing demand for scalable QPT schemes (Guo et al. 2019; Wu et al. 2021).
@@
Inspired by machine learning techniques, various methods have been proposed to alleviate the exponential scaling problem. Some statistical methods such as maximum-likelihood estimation (MLE) (Smolin, Gambetta, and Smith 2012), Bayesian estimation (Granade, Combes, and Cory 2016), and projected-least-squares (PLS) (Surawy-Stepney et al. 2022) techniques have been adapted for QPT. A major challenge in current applications is the experimental and computational cost of estimating large-scale systems. A class of methods partially address this challenge, including compressed-sensing QPT (CS-QPT) (Rodionov et al. 2014; Flammia et al. 2012), tensor network QPT (TN-QPT) (Torlai et al. 2023), but at the cost of assuming specific structures of the unknown quantum process. Recently, quantum machine learning combined with variational quantum circuits (VQCs) has received extensive attention (Abbas et al. 2021; Torlai et al. 2018). Another promising class of methods focuses on reconstructing a unitary quantum process by inverting the dynamics using variational algorithms (Xue et al. 2022; Carolan et al. 2020).
@@
Indeed, although these methods employ different techniques for reconstruction, they all involve a crucial step: query, which requires selecting a subset of quantum states from the informationally complete set as input of the unknown process. It is worth noting that quantum states with greater informativeness typically improving reconstruction more efficiently. However, previous works commonly use a straightforward strategy to select the subset randomly, overlooking differences in informativeness among quantum states. Since querying the quantum system requires multiple experiments that can be prohibitively costly, it is always the case that there are not enough quantum states for high-quality reconstruction, especially in large-scale quantum systems. Building on this observation, this work does not aim to replace these methods but to propose a general framework that further enhances them by optimizing the query step. Specifically, we focus on the informativeness of quantum states, aiming to select a subset of quantum states that improves the reconstruction most efficiently.
@@
AL is a promising machine learning paradigm that helps reduce the annotation costs by intelligently selecting a subset of informative samples for labeling (Settles 2009). With the aim of reducing experiment costs, AL has been extended from computer science to physics and applied successfully in various subfields such as condensed matter physics (Zhang et al. 2019; Yao et al. 2020), high-energy physics (Caron et al. 2019), and quantum information (Ding et al. 2020; Lange et al. 2023; Dutt et al. 2023). Inspired by this, we first utilize the differences in infromativeness among quantum states and propose a general framework for AL to adaptively select a subset of informative quantum states that improves the reconstruction most efficiently. In particular, we introduce a learning framework that leverages the widely-used VQCs to perform the QPT task and integrate our AL algorithms into the query step. Additionally, we design and evaluate three various types of AL algorithms with the aim of offering choices suited to different scenarios, including committee-based, uncertainty-based, and diversity-based, each exhibiting distinct advantages in terms of performance and computational cost (Kumar and Gupta 2020).
Contributions.
@@
The main contributions of this work are summarized as follows. First, we propose a general framework for AL that enhances previous methods by optimizing the crucial query step. To the best of our knowledge, this is the first framework that utilizes the differences in informativeness among quantum states. In particular, we introduce a learning framework that leverages the widely-used VQC to perform the QPT task and integrate our AL algorithms into the query step. Second, we design and evaluate three various types of AL algorithms: committee-based, uncertainty-based, and diversity-based, each exhibiting distinct advantages in terms of performance and computational cost. Third, numerical results demonstrate that our algorithms achieve significantly improved reconstruction compared to the baseline that selects a subset of quantum states randomly. Additionally, these results further indicate the importance of our algorithms, especially in large-scale quantum systems. Lastly, we provide a guideline for selecting algorithms suitable for different scenarios.
Preliminaries
Figure 1: Hybrid quantum-classical procedure of executing and optimizing a variational quantum circuit (VQC).
In this section, we briefly introduce the basic concepts of quantum computation and quantum information that are necessary in this work. Interested readers are recommended to the celebrated textbook by Nielsen and Chuang (Nielsen and Chuang 2001).
Qubit, Quantum State.
Qubit is the fundamental concept of quantum computation and quantum information. Unlike a classical bit is strictly binary, a qubit can be in a superposition of both binary states. A single-qubit quantum state is a linear combination of two basis states, represented as $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ in Dirac notation, where $\alpha,\beta\in\mathbb{C}$ and $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$. Here, $|0\rangle$ and $|1\rangle$ denote the basis states $\left[1,0\right]^{T}$ and $\left[0,1\right]^{T}\in\mathbb{C}^{2}$, respectively. Similarly, an $n$-qubit quantum state can be represented as $|\psi\rangle=\sum_{i=0}^{2^{n}-1}\alpha_{i}|i\rangle$, where $\alpha_{i}\in\mathbb{C}$, $|i\rangle\in\mathbb{C}^{2^{n}}$ for each $i$, and $\sum_{i}\left|\alpha_{i}\right|^{2}=1$.
Unitary Transformation, Pauli Matrices.
An $n$-qubit unitary transformation is defined as an operator $U\in\mathbb{C}^{2^{n}\times 2^{n}}$ that satisfies $U^{\dagger}U=I$, where $U^{\dagger}$ denotes the conjugate transpose of $U$. The operator $U$ transforms the quantum state of the system from $|\psi\rangle$ to $U|\psi\rangle$. Pauli matrices are key examples of such unitary transformations:
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad Y=\begin{pmatrix}0&-\textnormal{i}\\ \textnormal{i}&0\end{pmatrix},\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
Quantum Gate, Quantum Circuit.
Some important quantum gates featured extensively in this work include the Pauli rotations $\left\{R_{P}(\theta)=\textnormal{e}^{-\textnormal{i}\frac{\theta}{2}P}\mid P\in\left\{X,Y,Z\right\}\right\}$, which are exponential forms of the Pauli matrices. Additionally, the controlled-X (CNOT) gate, represented as $\text{CNOT}=|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes X$, is used to generate quantum entanglement among qubits. A sequence of quantum gates applied to a system of qubits forms a quantum circuit, at the end of which the qubits are measured to obtain classical values.
Figure 2: Framework of active learning with variational quantum circuits for quantum process tomography.
Variational Quantum Circuit.
VQC is a hybrid quantum-classical approach that leverages the strengths of both quantum and classical computation. It is a class of quantum circuits with tunable parameters, optimized for a specific purpose on a classical computer, while the variational circuit itself is executed on a quantum computer, as illustrated in Fig. 1. This approach has the advantage of keeping the quantum circuit depth shallow and hence mitigating noise. Consequently, VQCs have been developed for a wide range of applications, covering essentially all of the applications that researchers had envisaged for quantum computers, and they appear to be the best hope for obtaining quantum advantage (Cerezo et al. 2021).
Approximating Unitary Quantum Process with Variational Quantum Circuits
@@
In this section, we discuss how the design leverages widely-used VQCs to perform the QPT task and provide a detailed explanation of the learning framework.
Model
Unknown Quantum Process.
@@
Let $\mathcal{U}$ be an unknown unitary quantum process. It is reasonable to compile it into a variational quantum circuit (Khatri et al. 2019; Xue et al. 2022). More specifically, we can approximate $\mathcal{U}$ using a VQC denoted by $\mathcal{C}(\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ represents the parameters to be optimized. Consequently, all the information about $\mathcal{U}$ is encoded in the parameters $\boldsymbol{\theta}$, enabling us to systematically reconstruct $\mathcal{U}$ from these parameters using a classical computer.
Variational Quantum Circuit Ansatz.
@@
The ansatz of VQC is shown in Fig. 2. It is composed of interlaced layers using single-qubit Pauli rotation gates and two-qubit CNOT gates, designed to quickly generate entanglement between qubits. Thus, it makes it possible to approximate complicated quantum processes. Each two-qubit layer is counted as a depth and each single-qubit layer contains three rotational gates ($R_{z}$, $R_{y}$ and $R_{z}$) on each qubit. For such a circuit with $n$ qubits and $d$ depths, the total number of parameters is $3n(d+1)$.
Learning Framework
@@
The learning framework encompasses three key components: quantum state preparation, query, and training, as shown in Fig. 2.
Quantum State Preparation.
@@
A quantum state is prepared by applying a unitary transformation $U$ to the initial state: $|0\rangle^{\otimes n}\rightarrow U|0\rangle^{\otimes n}$. Specifically, the unitary transformation $U$ is selected from tensor products of single-qubit Pauli operators $\left\{I,X,Y,Z\right\}^{\otimes n}$. We generate an informationally complete set of quantum states, denoted as $D=\left\{\left|\psi_{1}\right\rangle,\left|\psi_{2}\right\rangle,\ldots,\left|\psi_{N}\right\rangle\right\}$, where $N=4^{n}$.
Query.
@@
We refer to the unknown unitary process $\mathcal{U}$ as an oracle that can be queried. When a quantum state $|\psi\rangle$ is queried to the oracle, it returns the state $|\psi^{\mathrm{ideal}}\rangle=\mathcal{U}|\psi\rangle$ as its true label. This procedure allows us to construct a training dataset.
Training.
@@
Given a training dataset $D_{l}=\left\{\left(\left|\psi_{1}\right\rangle,\left|\psi_{1}^{\mathrm{ideal}}\right\rangle\right),\left(\left|\psi_{2}\right\rangle,\left|\psi_{2}^{\mathrm{ideal}}\right\rangle\right),\ldots,\left(\left|\psi_{m}\right\rangle,\left|\psi_{m}^{\mathrm{ideal}}\right\rangle\right)\right\}$, the loss function is defined as the mean square error:
@@
$\mathcal{L}(\boldsymbol{\theta})$ $=\frac{1}{m}\sum_{i=1}^{m}\left\|\left|\psi_{i}(\boldsymbol{\theta})\right\rangle-\left|\psi_{i}^{\mathrm{ideal}}\right\rangle\right\|^{2}$
@@
$=\frac{1}{m}\sum_{i=1}^{m}\bigg{(}\left\langle\psi_{i}(\boldsymbol{\theta})\mid\psi_{i}(\boldsymbol{\theta})\right\rangle+\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}^{\mathrm{ideal}}\right\rangle$
@@
$\quad-\left\langle\psi_{i}(\boldsymbol{\theta})\mid\psi_{i}^{\mathrm{ideal}}\right\rangle-\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}(\boldsymbol{\theta})\right\rangle\bigg{)}$
@@
$=\frac{2}{m}\sum_{i=1}^{m}\left[1-\mathrm{Re}\left(\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}(\boldsymbol{\theta})\right\rangle\right)\right],$ (1)
where $\left|\psi_{i}\left(\boldsymbol{\theta}\right)\right\rangle=\mathcal{C}\left(\boldsymbol{\theta}\right)\left|\psi_{i}\right\rangle$ denotes the predicted label. The inner product on the right-hand side can be efficiently computed using a generalized SWAP-test algorithm on a quantum computer (Xue et al. 2022).
@@
To minimize the loss function, we employ a gradient-based optimization method to iteratively update the parameters $\boldsymbol{\theta}$. Using the chain rule, the gradient is computed as
$\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\theta_{i}}=\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}},$ (2)
@@
where $\mathcal{O}$ represents the observable corresponding to the generalized SWAP test. The first term, $\frac{\partial\mathcal{L}(\boldsymbol{\theta})}{\partial\mathcal{O}(\boldsymbol{\theta})}$, can be computed classically once the value of $\mathcal{O}(\boldsymbol{\theta})$ is obtained from the quantum computer. The second term, $\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}}$, is computed on the quantum computer using the parameter-shift rule (Mitarai et al. 2018):
$\frac{\partial\mathcal{O}(\boldsymbol{\theta})}{\partial\theta_{i}}=\frac{1}{2}\left[\mathcal{O}(\boldsymbol{\theta}_{i}^{+})-\mathcal{O}(\boldsymbol{\theta}_{i}^{-})\right],$ (3)
@@
where $\boldsymbol{\theta}_{i}^{\pm}$ denotes the parameter vector $\boldsymbol{\theta}$ with the $i$th parameter shifted by ±$\pm\frac{\pi}{2}$.
Method

### v2

# thanks: Corresponding authorthanks: Corresponding author

@@ Abstract
Quantum process tomography (QPT) is a fundamental tool for fully characterizing quantum systems. It relies on querying a set of quantum states as input to the quantum process. Previous QPT methods typically employ a straightforward strategy for randomly selecting quantum states, overlooking differences in informativeness among them. In this work, we propose a general active learning (AL) framework that adaptively selects the most informative subset of quantum states for reconstruction. We design and evaluate various AL algorithms and provide practical guidelines for selecting suitable methods in different scenarios. In particular, we introduce a learning framework that leverages the widely-used variational quantum circuits (VQCs) to perform the QPT task and integrate our AL algorithms into the query step. We demonstrate our algorithms by reconstructing the unitary quantum processes resulting from random quantum circuits with up to seven qubits. Numerical results show that our AL algorithms achieve significantly improved reconstruction, and the improvement increases with the size of the underlying quantum system. Our work opens new avenues for further advancing existing QPT methods.
preprint: APS/123-QED
I Introduction
@@
Quantum process tompography (QPT), used for reconstruction of an unknown quantum process from measurement data, is a fundamental tool for the diagnostic and full characterization of quantum systems [1]. Standard quantum process tomography (SQPT) relies on an informationally complete set of quantum states as input, which inevitably leads to an algorithmic complexity in terms of the number of measurements and classical post-processing that scales exponentially with the system size [2]. Due to these limitations, SQPT has only been experimentally implemented up to three qubits [3, 4]. However, with the rapid advancement of quantum computing hardware, there is a growing demand for scalable QPT schemes [5, 6].
@@
Inspired by machine learning techniques, various methods have been proposed to alleviate the exponential scaling problem. Some statistical methods such as maximum-likelihood estimation [7], Bayesian estimation [8], and projected-least-squares [9] techniques have been adapted for QPT. A major challenge in current applications is the experimental and computational cost of estimating large-scale systems. A class of methods partially address this challenge, including compressed-sensing QPT [10, 11], tensor network QPT [12], but at the cost of assuming specific structures of the unknown quantum process. In recent years, quantum machine learning combined with variational quantum circuits (VQCs) has received extensive attention [13, 14]. Another promising class of methods focuses on reconstructing a unitary quantum process by inverting the dynamics using variational algorithms [15, 16].
@@
Indeed, although these QPT methods employ different techniques for reconstruction, they all involve a crucial step: query, which requires selecting a subset of quantum states from the informationally complete set as inputs to the unknown process. It is worth noting that quantum states with greater informativeness typically improving reconstruction more efficiently. However, previous works typically use a straightforward strategy to select the subset randomly, overlooking differences in informativeness among quantum states. Since querying the quantum system requires multiple experiments that can be prohibitively costly, it is always the case that there are not enough quantum states for high-quality reconstruction, especially in large-scale quantum systems.
@@
Building on this observation, this work does not aim to replace existing QPT methods, but rather proposes a general framework that further enhances them by optimizing the crucial query step. Specifically, we focus on the informativeness of quantum states, aiming to select a subset of quantum states that improves the reconstruction most efficiently. Active learning (AL) is a promising machine learning paradigm that helps reduce the annotation costs by intelligently selecting a subset of informative samples for labeling [17]. With the aim of reducing experiment costs, AL has been extended from computer science to physics and applied successfully in various subfields, such as condensed matter physics [18, 19], high-energy physics [20], and quantum information [21, 22, 23].
@@
In this paper, we propose a general AL framework that adaptively selects the most informative subset of quantum states for reconstruction. To the best of our knowledge, this is the first framework that explicitly utilizes the differences in informativeness among quantum states for QPT. We design and evaluate three various types of AL algorithms with the aim of offering choices suited to different scenarios, including committee-based, uncertainty-based, and diversity-based, each exhibiting distinct advantages in terms of performance and computational cost [24]. In particular, we introduce a learning framework that leverages the widely-used VQCs to perform the QPT task and integrate our AL algorithms into the query step. Notably, we emphasize our AL framework is general as it can in principle be applied in the query step of any QPT method. We demonstrate our algorithms by reconstructing the unitary quantum processes resulting from random quantum circuits with up to seven qubits. Numerical results show that our AL algorithms achieve significantly improved reconstruction compared to the baseline method, which randomly selects a subset of quantum states. Moreover, the improvement increases with the size of the underlying quantum system.
This paper is organized as follows. In Sec. II, we introduce how widely-used VQCs are leveraged to perform the QPT task. We then provide a detailed explanation of the learning framework. In Sec. III, we propose a general AL framework that extends the learning framework by optimizing its query step through the exploitation of differences in informativeness among quantum states. We further design various AL algorithms. In Sec. IV, we demonstrate our algorithms with numerical simulations on the quantum processes of randomly generated circuits with up to seven qubits. We conclude in Sec. V.
II Learning Unitary Quantum Process with Variational Quantum Circuits
Figure 1: Schematic of our proposed active learning framework for quantum process tomography. The framework consists of three key components: Quantum State Preparation, Querying, and Training. In the Quantum State Preparation step, we prepare an informationally complete set of quantum states to initialize the unlabeled pool $D_{u}$. We show our active learning cycle as follows: it consists of two steps, i.e., Querying and Training. In the Querying step, we select the most informative quantum state $|\psi^{\star}\rangle$ via the designed acquisition function (including various types, e.g., Committee-Based, Uncertainty-Based, and Diversity-Based). We then query the unknown unitary process $\mathcal{U}$ to obtain its true label $\mathcal{U}|\psi^{\star}\rangle$. The newly labeled quantum state $(|\psi^{\star}\rangle,\mathcal{U}|\psi^{\star}\rangle)$ is subsequently added to the labeled pool $D_{l}$. In the Training step, we retrain the VQC model using the updated labeled pool $D_{l}$ to better approximate the unknown unitary process $\mathcal{U}$. This cycle repeats until the labeling budget is exhausted or the VQC model achieves satisfactory performance.
II.1 Model
@@
In this section, we discuss how to leverage widely-used VQCs to perform the QPT task and provide a detailed explanation of the learning framework.
@@
Unknown Quantum Process.— Let $\mathcal{U}$ be an unknown unitary quantum process. It is reasonable to compile it into a VQC [25, 16]. More specifically, we can approximate $\mathcal{U}$ using a VQC denoted by $\mathcal{C}(\bm{\theta})$, where $\bm{\theta}$ represents the parameters to be optimized. Consequently, all the information about $\mathcal{U}$ is encoded in the parameters $\bm{\theta}$, enabling us to systematically reconstruct $\mathcal{U}$ from these parameters using a classical computer.
@@
Variational Quantum Circuit Ansatz.— The ansatz of VQC is shown in Fig. 1. It is composed of interlaced layers using single-qubit Pauli rotation gates and two-qubit CNOT gates. This structure is designed to quickly generate entanglement between qubits, thus making it possible to approximate complicated quantum processes. Each two-qubit layer is counted as a depth and each single-qubit layer contains three rotational gates ($R_{z}$, $R_{y}$, and $R_{z}$) on each qubit, where
$R_{y}(\theta)=\begin{pmatrix}\cos\frac{\theta}{2}&-\sin\frac{\theta}{2}\\ \sin\frac{\theta}{2}&\cos\frac{\theta}{2}\end{pmatrix},\quad R_{z}(\theta)=\begin{pmatrix}\textnormal{e}^{-\textnormal{i}\frac{\theta}{2}}&0\\ 0&\textnormal{e}^{\textnormal{i}\frac{\theta}{2}}\end{pmatrix}.$
The combination of $R_{z}$, $R_{y}$, and $R_{z}$ rotations can realize any single-qubit rotation by adjusting the parameters. For such an $n$-qubit circuit with depth $k$, the total number of parameters is $3n(k+1)$.
II.2 Learning Framework
@@
The learning framework encompasses three key components: quantum state preparation, querying, and training, as shown in Fig. 1.
@@
Quantum State Preparation.— A quantum state is prepared by applying a unitary transformation $U$ to the initial state: $|0\rangle^{\otimes n}\rightarrow U|0\rangle^{\otimes n}$. Specifically, the unitary transformation $U$ is selected from tensor products of single-qubit Pauli operators $\left\{I,X,Y,Z\right\}^{\otimes n}$. We generate an informationally complete set of quantum states, denoted as $D=\left\{\left|\psi_{1}\right\rangle,\left|\psi_{2}\right\rangle,\ldots,\left|\psi_{N}\right\rangle\right\}$, where $N=4^{n}$.
@@
Querying.— We refer to the unknown unitary process $\mathcal{U}$ as an oracle that can be queried. When a quantum state $|\psi\rangle$ is queried to the oracle, it returns the state $|\psi^{\mathrm{ideal}}\rangle=\mathcal{U}|\psi\rangle$ as its true label. This procedure allows us to construct a training dataset.
@@
Training.— Given a training dataset $D_{l}=\left\{\left(\left|\psi_{1}\right\rangle,\left|\psi_{1}^{\mathrm{ideal}}\right\rangle\right),\left(\left|\psi_{2}\right\rangle,\left|\psi_{2}^{\mathrm{ideal}}\right\rangle\right),\ldots,\left(\left|\psi_{m}\right\rangle,\left|\psi_{m}^{\mathrm{ideal}}\right\rangle\right)\right\}$, the loss function is defined as the mean square error:
@@
$\mathcal{L}(\bm{\theta})$ $=\frac{1}{m}\sum_{i=1}^{m}\left\|\left|\psi_{i}(\bm{\theta})\right\rangle-\left|\psi_{i}^{\mathrm{ideal}}\right\rangle\right\|^{2}$
@@
$=\frac{1}{m}\sum_{i=1}^{m}\bigg{(}\left\langle\psi_{i}(\bm{\theta})\mid\psi_{i}(\bm{\theta})\right\rangle+\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}^{\mathrm{ideal}}\right\rangle$
@@
$\quad-\left\langle\psi_{i}(\bm{\theta})\mid\psi_{i}^{\mathrm{ideal}}\right\rangle-\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}(\bm{\theta})\right\rangle\bigg{)}$
@@
$=\frac{2}{m}\sum_{i=1}^{m}\left[1-\mathrm{Re}\left(\left\langle\psi_{i}^{\mathrm{ideal}}\mid\psi_{i}(\bm{\theta})\right\rangle\right)\right],$ (1)
[…]
@@
To minimize the loss function, we employ a gradient-based optimization method to iteratively update the parameters $\bm{\theta}$. Using the chain rule, the gradient is computed as
$\frac{\partial\mathcal{L}(\bm{\theta})}{\partial\theta_{i}}=\frac{\partial\mathcal{L}(\bm{\theta})}{\partial\mathcal{O}(\bm{\theta})}\frac{\partial\mathcal{O}(\bm{\theta})}{\partial\theta_{i}},$ (2)
@@
where $\mathcal{O}$ represents the observable corresponding to the generalized SWAP test. The first term, $\frac{\partial\mathcal{L}(\bm{\theta})}{\partial\mathcal{O}(\bm{\theta})}$, can be computed classically once the value of $\mathcal{O}(\bm{\theta})$ is obtained from the quantum computer. The second term, $\frac{\partial\mathcal{O}(\bm{\theta})}{\partial\theta_{i}}$, is computed on the quantum computer using the parameter-shift rule [26]:
$\frac{\partial\mathcal{O}(\bm{\theta})}{\partial\theta_{i}}=\frac{1}{2}\left[\mathcal{O}(\bm{\theta}_{i}^{+})-\mathcal{O}(\bm{\theta}_{i}^{-})\right],$ (3)
@@
where $\bm{\theta}_{i}^{\pm}$ denotes the parameter vector $\bm{\theta}$ with the $i$-th parameter shifted by ±$\pm\frac{\pi}{2}$.
III Method